# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

### v1

@@ 5 Future Work
[set/.style = draw]
[…]
“Die Dinge wurden kompliziert”Score:1
“Die Dinge erhalten kompliziert”Score:0.5;
[…]
In German, the verb ”bekommen” means ”to get” or
”to receive,” but in this context, it doesn’t convey the
intended meaning effectively. Instead, we can use the
verb ”werden” which means ”to become” to better
capture the sense of change or transformation.;
[…]
“became” Source & “werden” Hypothesis
Score;
[…]
Importance((“became”,“bekommen”)): 0.1
Importance((“Things”,“Die Dinge”)): 0.6;
[…]
Neuron 1423: bekommen
Neuron 5234: werden;
[…]
Explanation Types;
[…]
Ground Truth: “Things became complicated”
Hypothesis: “Die Dinge bekamen kompliziert”
Score: ;
[[sep=6.2pt]-Computer Modern Rightarrow[length=3mm, width=3mm], line width=0.7mm, shorten ¡=8pt, shorten ¿=8pt] (A) edge (F) (B) edge (F) (C) edge (F) (D) edge (F) (E) edge (F) ;
@@
Exploring present explanation types Based on our taxonomy (see Figure 4), we identify unexplored areas of explainable MT evaluation (empty table cells). We find that only feature importance techniques are thoroughly explored on all levels of model access. The other methods mostly explain in black box settings or are based on metrics that are explainable by design. We suppose that gradients, attention weights and embeddings have so far mostly been applied in the field of feature importance, as there is often already a 1-1 relationship their values have towards input features. Also, due to the two recent shared tasks, feature importance has likely received the most attention.
Future work could consider enhancing quality aspect explanations — Row 1 in Figure 4 — by incorporating more model access. For example, a separate fluency score might be realizable by employing measurements on target side embeddings in unsupervised matching metrics. Current model-understanding techniques for MT metrics — Row 4-5 in Figure 4 — are mostly applied in a black box manner. For these, using more model access might also provide further insights. For example, analyses of the contextualized embedding space and how it affects a metric score could explain the metric on a more general level. Many works consider the probing of transformer-based architectures (e.g. Tenney et al. 2019; Liu et al. 2019; Sajjad et al. 2022b) and it would be interesting to see if their results can inform the metric development.
@@
Exploring other explanation types Besides the works in our taxonomy, other fields of NLP and explainable AI in general have proposed further explanation types, e.g. counterfactuals, natural language explanations and rule extraction (e.g. Madsen et al. 2022). We illustrate these options with an example in Figure 6. We already introduced counterfactuals in §2.2, stating their role in evaluations of robustness (as adversarial attacks) and fairness. It remains to explore if single sentence explanations also carry explanatory value. In the examples for counterfactuals in Figure 6, the original word “bekommen” is switched with the words “wurden” (to get) resp. “erhalten” (to receive) in two counterfactuals. Based on the scores achieved by these counterfactuals, a user could infer metric weaknesses. One technique to generate counterfactuals is for example described by Wu et al. (2021).
@@
Natural language explanations are explanations in free text form (e.g. Wiegreffe and Marasovic 2021). This is an explainability type where the generative abilities of recent large language models (LLM), like ChatGPT and GPT4, can be leveraged. The works by Fu et al. (2023); Lu et al. (2023); Kocmi and Federmann (2023) already explore the use of LLMs as metrics and extract the metric scores from LLM outputs (or generation probabilities) that contain (or could easily be extended to provide) natural language descriptions of why a certain score is assigned. In contrast, Xu et al. (2023) explicitly highlight the explainability aspect of their model. In our taxonomy, we grouped them under fine-grained error explanations, as this is their main goal. However, they also produce natural language explanations for each error.151515We again note that these techniques will not necessarily describe a metric’s internal workings, but might be usable due to providing plausible explanations, see §2.2. In these works, the explanations are rather used as a byproduct, while an evaluation of the explanation quality and further use cases remains for future work.
@@
The example in Figure 6 was generated with the help of ChatGPT, which shows that such an approach might yield plausible explanations. An example from other NLP domains is Rajani et al. (2019), who generate explanations for commonsense reasoning. In NLG in general, many other approaches for LLM-based metrics have been constructed since the release of ChatGPT. These could potentially be extended to MT metrics and explainability (e.g. Chiang and yi Lee 2023; Wang et al. 2023; Ji et al. 2023a; Chen et al. 2023; Liu et al. 2023).
@@
Recent works, describe prompting strategies like chain-of-thought (Wei et al. 2022) and tree-of-thought (Yao et al. 2023) where language models show performance improvements if they are giving explanations of their output. It could be an interesting research direction, to see if the explainability techniques summarized in this paper can improve metric performance in this way, or if these techniques can provide reasonable explanations (that are at least plausible).
@@
Neuron Explanations are another potential use case of LLMs to separately explain each Neuron in the underlying networks. Bills et al. (2023) recently showed that GPT4 can (to some extent) summarize highlighted neuron activation in textual input of GPT2 to the same extent humans can. This allows to successfully simulate the behavior and effects of some neurons. For MT metrics, this would allow to determine the role certain neurons play to determine scores when specific features like “became is used” are present in the metric input. At the moment, these methods however still have some limitations Bills et al. (2023).
@@
Interactive Explanations Due to their context size, instruction fine-tuning and emergent abilities of recent LLMs allow to interactively ask questions and adapt based on the answers (e.g. OpenAI 2023b). Jacovi et al. (2023) state that interactivity is a key point of successfully communicating the explained subject to an explainee. LLM-based explanations (like natural language/neuron explanations …) might achieve this. For example, a user could ask for more detailed information for a mistranslation that was noted in a first explanation.
@@
New approaches could furth consider displaying multiple explanation types together, like Xu et al. (2023) who present fine-grained error annotations with natural language explanations.
@@
Future evaluation and concepts Current work scarcely performs human evaluation of explanations. This is partly because the focus is often not on the explanations but other aspects, like the sentence-level scores. One form of human evaluation (for plausibility) is often to use pre-computed human explanations and check the correlation to them (e.g. Zerva et al. 2022). Current works on MT metrics do, however, not evaluate whether the explanations are actually helpful to explain the concepts they describe to their audience. This means, whether the explanations manage to help the explainees to adapt their own theory of mind to be a coherent abstraction of the real problem Jacovi et al. (2023). Hence, we encourage future work to address this, e.g. by considering simulatability Hase and Bansal (2020). Regarding plausibility, it could also be interesting to evaluate whether plausible explanations of incorrect metric scores can actually deceive non-experts or experts (with the current state-of-the-art), as warned by Jacovi and Goldberg (2021); Jin et al. (2023).
@@
As noted throughout the paper, the degree of faithfulness an explanation has to the metric’s internal decision process is an important property (Jacovi and Goldberg 2020), especially for metric debugging. Current works on decision-understanding of MT metrics, however, mostly dismiss faithfulness. Future work might consider the evaluation of metric faithfulness, to enable these use-cases. We note that this evaluation will be complex, due to faithfulness being multi-faceted (Jacovi and Goldberg 2020). For example, lower-granularity scores can faithfully explain sentence-level scores that are aggregated from them, however this does not explain how the actual black box component produced the lower-granularity scores. Further, there are pitfalls that exist in faithfulness evaluation (Ju et al. 2022), warning against the usage of common methods like area over the perturbation curve scores (AOPC) (Samek et al. 2017). Therefore, future work should first consider what constitutes faithful explanations in terms of metrics and design evaluation approaches.
@@
Further, we note that current works mostly consider explaining sentence-level scores. Future work, might also consider other metric granularity types more.
@@
Exploring domain boundaries Future work might also consider under-explored use cases of explainability for MT metrics. Currently, only one work considers the social biases of MT metrics Sun et al. (2022) and none considers the usage of explainability for semi-automatic labeling of training data. Besides examining just the metrics, future work could also inspect the relation of explainable MT metrics, especially plausible-only explanations, to explainable MT (e.g. Stahlberg et al. 2018), as explanations of both will probably be similar. Lastly, future work could explore providing similar taxonomies as ours that are comprising the whole field of explainable natural language generation metrics.
@@ 6 Related Work
A large number of conceptual works and surveys have been conducted in the area of explainability (e.g. Lipton 2016; Biran and Cotton 2017; Adadi and Berrada 2018; Doshi-Velez and Kim 2017; Došilović et al. 2018; Guidotti et al. 2018; Arya et al. 2019; Carvalho et al. 2019; Gilpin et al. 2019; Miller 2019; Barredo Arrieta et al. 2020; Bodria et al. 2023; Linardatos et al. 2021).
@@
General surveys on explainable NLP are related to our work, as they highlight how explainability techniques have been used in NLP and might further guide the development of future approaches for explainable MT metrics (e.g. Vijayakumar 2023; Sajjad et al. 2022a; Gurrapu et al. 2023; Lertvittayakumjorn and Toni 2021; Saha et al. 2022; Sun et al. 2021; Zini and Awad 2022; Qian et al. 2021; Madsen et al. 2022). Besides these works, there are also surveys of robustness (e.g. Goyal et al. 2023) and fairness (e.g. Blodgett et al. 2020), which are related, as discussed in §2.2.
@@ 7 Conclusion
In this work, we discuss audiences, goals and properties for explainable machine translation metrics, a nascent field that may help further overcome the dominance of classical low-quality evaluation metrics. We also survey and categorize recent approaches on explainable MT metrics into a taxonomy, highlighting their results and limitations. Currently, two dominant approaches to explainability for MT metrics are (1) feature importance explanations that highlight erroneous words in source and hypothesis to explain sentence-level scores and (2) perturbation robustness approaches that check a metric’s robustness to manually devised types of input perturbations. We also identify a current trend towards (3) fine-grained error explanations likely caused by the recent improvements of LLMs like ChatGPT and by the adaptation of MQM annotations in the WMT metrics shared tasks. A major weakness of the current realization of (1) is that error highlights do not consider the correspondence between words in source and target. Further, a weakness of (2) is that perturbation types have to be manually defined, an issue that might be tackled by using more general approaches of adversarial attacks in the future (we discuss this issue in more detail in an earlier version of this paper; Leiter et al. 2022a). Generally, most current evaluations of explainability approaches for MT metrics do not consider the faithfulness of explanations, limiting their use in metric debugging and high-risk scenarios. We also present a vision of next-generation approaches to explainable evaluation metrics, which should help fix the problems of the above paradigms and provide guidance to explore unexplored areas. Here, we also urge future work to consider the desiderata and implications of faithfulness for explainable MT metrics.
@@
Acknowledgements
Since November 2022 Christoph Leiter is financed by the BMBF project “Metrics4NLG”. Piyawat Lertvittayakumjorn had been financially supported by Anandamahidol Foundation, Thailand, from 2015-2021. He mainly contributed to this work only until September 2022 while affiliated with Imperial College London (before joining Google as a research scientist after that). Marina Fomicheva mainly contributed to this work until April 2022. Wei Zhao is supported by the Klaus Tschira Foundation, Heidelberg, Germany. Yang Gao mainly contributed to this work before he joined Google Research in December 2021. Steffen Eger is financed by DFG Heisenberg grant EG 375/5–1 and by the BMBF propject “Metrics4NLG”.

### v2

@@ 5 Future Work
Exploring present explanation types Based on our taxonomy (see Figure 3), we identify unexplored areas of explainable MT evaluation (empty table cells). We find that only feature importance techniques are thoroughly explored on all levels of model access. The other methods mostly explain in black box settings or are based on metrics that are explainable by design. We suppose that gradients, attention weights, and embeddings have, so far, mostly been applied in the field of feature importance, as there is often already a one-to-one relationship between their values and input features. Also, due to the recent shared tasks, feature importance has likely received the most attention.
Future work could consider enhancing quality aspect explanations — Row 1 in Figure 3 — by incorporating more model access. For example, a separate fluency score might be realizable by employing measurements on target side embeddings in unsupervised matching metrics. Current model-understanding techniques for MT metrics — Row 4-5 in Figure 3 — are mostly applied in a black box manner. For these, using more model access might also provide further insights. For example, analyses of the contextualized embedding space and how it affects a metric score could explain the metric on a more general level. Many works consider the probing of transformer-based architectures (e.g., Tenney et al., 2019; Liu et al., 2019; Sajjad et al., 2022b) and it would be interesting to see if their results can inform the metric development.
@@
Exploring other explanation types Besides the works in our taxonomy, other fields of NLP and explainable AI in general have proposed further explanation types, e.g., counterfactuals, natural language explanations and rule extraction (e.g., Madsen et al., 2022). We illustrate these options with an example in Figure 5. We already introduced counterfactuals in §2.2, stating their role in evaluations of robustness (as adversarial attacks) and fairness. It remains to explore if single sentence explanations also carry explanatory value. In the examples for counterfactuals in Figure 5, the original word “bekommen” is switched with the words “wurden” (to get) resp. “erhalten” (to receive) in two counterfactuals. Based on the scores achieved by these counterfactuals, a user could infer metric weaknesses. One technique to generate counterfactuals is for example described by Wu et al. (2021).
@@
Natural language explanations are explanations in free text form (e.g., Wiegreffe and Marasovic, 2021). This is an explainability type where the generative abilities of recent large language models (LLM), like ChatGPT and GPT4, can be leveraged. The works by Fu et al. (2023), Lu et al. (2023b), Kocmi and Federmann (2023b), Kocmi and Federmann (2023a), Fernandes et al. (2023) and Xu et al. (2023) already explore the use of LLMs as metrics and extract the metric scores from LLM outputs (or generation probabilities) that contain (or could easily be extended to provide) natural language descriptions of why a certain score is assigned. In our taxonomy, we grouped them under fine-grained error explanations, as this is their main goal.151515We again note that these techniques will not necessarily describe a metric’s internal workings, but might be helpful due to providing plausible explanations, see §2.2. In these works, the explanations are mostly used as a byproduct, while a thorough evaluation of the explanation quality and further use cases remains for future work.
The example in Figure 5 was generated with the help of ChatGPT, which shows that such an approach might yield plausible explanations.161616We used ChatGPT in June 2023 with following prompts and asked to rephrase the output once:
Ground Truth: “Things got complicate”
Hypothesis: “Die Dinge bekamen kompliziert”
Is this translation of the ground truth correct? Why or why not?
An example from other NLP domains is Rajani et al. (2019), who generate explanations for commonsense reasoning.
@@
In NLG in general, many other approaches for LLM-based metrics have been constructed since the release of ChatGPT. These could potentially be extended to MT metrics and explainability (e.g., Chiang and Lee, 2023; Wang et al., 2023; Ji et al., 2023; Chen et al., 2023; Liu et al., 2023).
As for example Kocmi and Federmann (2023a) point out, it it dangerous for academia to rely on proprietary models, especially for evaluation purposes. This includes the following issues: Metrics based on proprietary LLMs (1) could be influenced by the LLM providers, (2) could change or become unavailable at any time, (3) could carry more unexpected biases due to unknown training data and (4) could be contaminated in their evaluation by being trained on common test sets.
@@
Recent works, describe prompting strategies like chain-of-thought (Wei et al., 2022) and tree-of-thought (Yao et al., 2023) where language models show performance improvements if they are giving explanations of their output. While some of the works that are described in our taxonomy already employ such techniques to improve metrics a systematic exploration of the effect on metric scores, or whether these techniques can provide reasonable explanations (that are at least plausible), remains. A first work that considers this systematic exploration is the EVAL4NLP 2023 shared task (Leiter et al., 2023) where participants should improve metrics scores only by prompting. Their results on a small scale experiment indicate that the participants’ generated natural language explanations are not helpful to humans, in most cases.
@@
Neuron Explanations are another potential use case of LLMs to separately explain each Neuron in the underlying networks. Bills et al. (2023) recently showed that GPT4 can (to some extent) summarize highlighted neuron activation in textual input of GPT2 to the same extent humans can. This allows to successfully simulate the behavior and effects of some neurons. Interestingly, they also find error related neurons, like a “post-typo-neuron”. For MT metrics, this would allow to determine the role certain neurons play to determine scores when specific features like “became is used” are present in the metric input. At the moment, the success of neuron explanation methods is still limited (Bills et al., 2023).
@@
Interactive Explanations: Due to their context size, instruction fine-tuning and emergent abilities of recent LLMs allow to interactively ask questions and adapt based on the answers (e.g., OpenAI, 2023a). Jacovi et al. (2023) state that interactivity is a key point of successfully communicating the explained subject to an explainee. LLM-based explanations (like natural language/neuron explanations …) might achieve this. For example, a user could ask for more detailed information for a mistranslation that was noted in a first explanation.
@@
New approaches could further consider displaying multiple explanation types together, like Xu et al. (2023) who present fine-grained error annotations with natural language explanations.
@@
Future evaluation and concepts Current work scarcely performs human evaluation of explanations. This is partly because the focus is often not on the explanations but other aspects, like the sentence-level scores. One form of human evaluation (for plausibility) is often to use pre-computed human explanations and check the correlation to them (e.g., Zerva et al., 2022). Current works on MT metrics do, however, not evaluate whether the explanations are actually helpful to explain the concepts they describe to their audience. This means, whether the explanations manage to help the explainees to adapt their own theory of mind to be a coherent abstraction of the real problem Jacovi et al. (2023). Hence, we encourage future work to address this, e.g., by considering simulatability Hase and Bansal (2020). Human evaluation may also be performed in downstream tasks, by checking performance with and without explanations. Regarding plausibility, it could also be interesting to evaluate whether plausible explanations of incorrect metric scores can actually deceive non-experts or even translation experts, as warned by Jacovi and Goldberg (2021) and Jin et al. (2023) (for general explainability use cases).
@@
As noted throughout the paper, the degree of faithfulness an explanation has to the metric’s internal decision process is an important property (Jacovi and Goldberg, 2020), especially for metric debugging and bias detection. Current works on decision-understanding of MT metrics, however, mostly dismiss faithfulness. Future work might consider the evaluation of metric faithfulness, to enable these use-cases. We note that this evaluation will be complex, due to faithfulness being multi-faceted (Jacovi and Goldberg, 2020). For example, lower-granularity scores can faithfully explain sentence-level scores that are aggregated from them, however this does not explain how the actual black box component produced the lower-granularity scores. Further, there are evaluation pitfalls that exist in faithfulness evaluation (Ju et al., 2022), ruling out common methods like area under the perturbation curve scores (AOPC). Therefore, future work should first consider what constitutes faithful explanations in terms of metrics and design evaluation approaches.
@@
Further, we note that current works mostly consider explaining sentence-level scores. Future work, might also consider other metric granularity types more, such as word-level and document-level metrics. Additionally, future work could explore the influence of language resource availability on the explanations (although it may largely depend on metric quality) and whether some explainable metric setups are not applicable in low-resource settings.
@@
Exploring domain boundaries Future work might also consider under-explored use cases of explainability for MT metrics. Currently, only one work considers the social biases of MT metrics Sun et al. (2022) and none considers the usage of explainability for semi-automatic labeling of training data. Besides examining just the metrics, future work could also inspect the relation of explainable MT metrics, especially plausible-only explanations, to explainable MT (e.g., Stahlberg et al., 2018), as explanations of both will probably be similar. Lastly, future work could explore providing similar taxonomies as ours that are comprising the whole field of explainable natural language generation metrics.
@@ 6 Related Work
A large number of conceptual works and surveys have been conducted in the area of explainability (e.g., Lipton, 2016; Biran and Cotton, 2017; Adadi and Berrada, 2018; Doshi-Velez and Kim, 2017; Došilović et al., 2018; Guidotti et al., 2018; Arya et al., 2019; Carvalho et al., 2019; Gilpin et al., 2018; Miller, 2019; Barredo Arrieta et al., 2020; Bodria et al., 2023; Linardatos et al., 2021).
@@
General surveys on explainable NLP are related to our work, as they highlight how explainability techniques have been used in NLP and might further guide the development of future approaches for explainable MT metrics (e.g., Vijayakumar, 2023; Sajjad et al., 2022a; Gurrapu et al., 2023; Lertvittayakumjorn and Toni, 2021; Saha et al., 2022; Sun et al., 2021; Zini and Awad, 2022; Qian et al., 2021; Madsen et al., 2022). Besides these works, there are also surveys of robustness (e.g., Goyal et al., 2023) and fairness (e.g., Blodgett et al., 2020), which are related, as discussed in §2.2.
@@ 7 Conclusion
In this work, we discuss audiences, goals and properties for explainable machine translation metrics, a nascent field that may help further overcome the dominance of classical low-quality evaluation metrics. We also survey and categorize recent approaches on explainable MT metrics into a taxonomy, highlighting their results and limitations. Currently, two dominant approaches to explainability for MT metrics are (1) feature importance explanations that highlight erroneous words in source and hypothesis to explain sentence-level scores and (2) perturbation robustness approaches that check a metric’s robustness to manually devised types of input perturbations. We also identify a current trend towards (3) fine-grained error explanations likely caused by the recent improvements of LLMs like ChatGPT and by the adaptation of MQM annotations in the WMT metrics shared tasks. A major weakness of the current realization of (1) is that error highlights do not consider the correspondence between words in source and target. Further, a weakness of (2) is that perturbation types have to be manually defined, an issue that might be tackled by using more general approaches of adversarial attacks in the future (we discuss this issue in more detail in an earlier version of this paper; Leiter et al., 2022a). Generally, most current evaluations of explainability approaches for MT metrics do not consider the faithfulness of explanations, limiting their use in metric debugging and high-risk scenarios. We also present a vision of future approaches to explainable evaluation metrics, which should help fix the problems of the above paradigms and provide guidance to explore unexplored areas. Here, we also urge future work to consider the desiderata and implications of faithfulness for explainable MT metrics.
@@
Acknowledgments and Disclosure of Funding
Since November 2022 Christoph Leiter is financed by the BMBF project “Metrics4NLG”. Piyawat Lertvittayakumjorn had been financially supported by Anandamahidol Foundation, Thailand, from 2015-2021. He mainly contributed to this work only until September 2022 while affiliated with Imperial College London (before joining Google as a research scientist after that). Marina Fomicheva mainly contributed to this work until April 2022. Wei Zhao was supported by the Klaus Tschira Foundation and Young Marsilius Fellowship, Heidelberg, until December 2023. Yang Gao mainly contributed to this work before he joined Google Research in December 2021. Steffen Eger is financed by DFG Heisenberg grant EG 375/5–1 and by the BMBF propject “Metrics4NLG”.